A part of $6000 was invested at 6% annual interst and remaining at 7% annual interest. At the end of the year the total amount recieved was $6395. How much money was invested at each rate?
Answer by richwmiller(17219)

x+y=6000,
0.06*x+0.07*y=395
x=6000-y
0.06*(6000-y)+0.07*y=395
360-0.06y+0.07*y=395
0.01*y=35
y=3500 at 7%
x=6000-y
x=2500 at 6%
check
0.06*2500+0.07*3500=395
150+245=395
395=395
ok